### v1

# Large temperature-up-jump simulations of a binary Lennard-Jones system

Abstract
This paper presents simulations of the physical aging of a binary Kob-Andersen-type Lennard-Jones liquid following large temperature up-jumps from equilibrated states of high relaxation time. The purpose is to investigate how well the Tool-Narayanaswamy (TN) material-time concept works for this rather extreme case of aging. First the triangular relation of the potential energy is studied. This is found to be well obeyed, making it possible to define a potential-energy-based material time $\xi$. We proceed to study aging toward equilibrium at the final temperature 0.48 for jumps from the temperatures 0.43 and 0.37, monitoring the following five quantities: the potential energy, the self-intermediate scattering function, the mean-square displacement, the dynamic susceptibility $\chi_{4}$, and the non-Gaussian parameter $\alpha_{2}$. The TN material-time prediction is that all time-autocorrelation functions should collapse to only depend on the material-time difference $\xi_{2}-\xi_{1}$. This is found to work much better for the $0.43\to 0.48$ temperature jump than for the $0.37\to 0.48$ jump. Our findings thus confirm the general understanding that the TN aging formalism works best for systems that are never very far from equilibrium. This raises two questions for future work: Is the collapse significantly improved if each aging quantity is allowed its own material time? Can better collapse be obtained if the material-time is generalized to be defined locally in order to reflect dynamic heterogeneity?
I Introduction
It has been recognized for many years that physical aging of glasses is generally well described by the so-called material time [1, 2, 3, 4, 5, 6, 7]. This concept was introduced in 1971 by Ford Motor Co. engineer Narayanaswamy for optimizing the production of windscreens [8]. It is now known that the material-time idea rationalizes aging data of non-crystalline materials as different as oxide glasses [2], chalcogenide glasses [9], polymers [10, 11, 12, 13], and metallic glasses [14, 15]. The resulting so-called Tool-Narayanaswamy (TN) or TNM (Tool-Narayanaswamy-Moynihan) formalism has been used routinely for many years in industry in connection with glass production, as well as for predicting the long-time aging taking place in subsequent uses of glasses [2, 16].
Both production and applications of glasses involve complex temperature protocols. The conceptually simplest cases of physical aging are rapid jumps from a state of equilibrium at one temperature to a final temperature that is subsequently kept constant. Ideally, the “annealing” temperature is stabilized throughout the sample before any significant relaxation has taken place. Following a temperature jump, one or more properties of the system are monitored continuously as it slowly equilibrates, compare Fig. 1(a) that illustrates schematically aging following a temperature up jump. Due to the slow rate of heat conduction, such ideal equilibrium-to-equilibrium jumps are difficult to realize in experiments (though not impossible [3, 6]). In simulations, on the other hand, temperature jumps are easily implemented; here the challenge is to obtain the very long simulations needed to probe the relaxing region. If a global quantity is monitored, it is also difficult to get good statistics. The standard procedure of improving statistics by time averaging cannot be used due to the lack of time-translational invariance during aging.
In both simulations and experiments it is universally observed that temperature up and down jumps result in quite different responses. A temperature-down jump equilibrates faster and is more “stretched” in time than an up jump to the same final temperature; the latter is “compressed” in time when it eventually picks up speed. This difference is referred to as the “asymmetry of approach” [17, 13, 18, 19], with down and up jumps termed “self-retarded” and “self-catalyzed”, respectively [12]. Figure 1(b) gives examples of data for the potential energy following a single down jump and several up jumps to the same final temperature. The notable asymmetry between up and down jumps means that aging – from a mathematical point of view – is highly nonlinear. This reflects the long known fact [20] that during a temperature down jump, the structure gradually involves states of lower and lower energy, resulting in the dynamics slowing down, while the opposite happens for an up jump.
The TN formalism goes significantly beyond this qualitative reasoning known for almost a century [2, 13] by predicting that a material time $\xi=\xi(t)$ exists such that all temperature jumps to the same temperature are given by the same function of $\xi$ for the quantity monitored. This implies that if (laboratory) time is replaced by material time, physical aging is described as a standard linear-response convolution integral over the temperature-change history [8, 2, 21]. The conversion of the highly nonlinear aging phenomenon into a linear one constituted Narayanaswamy’s 1971 conceptual breakthrough [8].
Figure 1: Physical aging. (a) Schematic drawing of a temperature up jump starting and ending in thermal equilibrium (with time on the x-axis). Even though not all quantities relax in identical manner, the standard TN aging formalism operates with a single “global” clock controlling all relaxations toward equilibrium. (b) Results for the potential energy per particle following one down jump and several up jumps to the same final temperature $0.48$, illustrating the noted “asymmetry of approach” seen in all aging simulations and experiments. The black filled rectangles mark the potential energy at $T=0.48$.
The current paper presents a numerical investigation of the Kob-Andersen binary Lennard-Jones (LJ) mixture (slightly modified), following large temperature up jumps from states of equilibrium. The purpose is to investigate the limits of the TN material-time description for such jumps for which the TN formalism is more challenged than for down jumps. We define below the material time $\xi$ from the potential-energy time-autocorrelation function and investigate to which degree each of five different time-correlation functions during the aging collapse to be unique functions of the material-time difference $\xi_{2}-\xi_{1}$. We first discuss briefly the system simulated, as well as the triangular relation of Cugliandolo and Kurchan [22] and its connection to the material time.
II System simulated
We simulated a binary Lennard-Jones (mBLJ) model, which has the same parameters as and physics very similar to the standard Kob-Andersen model [23] but is at least 100 times less prone to crystallization [24]. The model is a mixture of 80% large particles (A) and 20% small particles (B) interacting by Lennard-Jones pair potentials, $v(r)=4\varepsilon((r/\sigma_{\alpha\beta})^{-12}-(r/\sigma_{\alpha\beta})^{-6})$ ($\alpha,\beta\in\{A,B\}$), in which $\sigma_{AA}=1$, $\sigma_{AB}=0.8$, $\sigma_{BB}=0.88$, $\varepsilon_{AA}=1$, $\varepsilon_{AB}=1.5$, $\varepsilon_{BB}=0.5$. The difference to the standard Kob-Andersen model is that instead of using shifted-potential cutoffs at relatively long distances, a shifted-force cutoff at 1.5 is used for the AA and BB interactions. This considerably weakens the like-particle attractions and thereby impedes crystallization by phase separation into a pure A phase [25]. The AB interaction also has a shifted-force cutoff, but at 2.5 [23]. All simulations were carried out with $N_{A}=6400$ A particles and $N_{B}=1600$ B particles. The quantities evaluated refer exclusively to the A particles because they dominate the structural relaxation. The simulations utilized the GPU-optimized MD code RUMD [26].
III Triangular relation
In developments two decades after Narayanaswamy that were entirely independent of the experimentally motivated TN formalism, Cugliandolo and Kurchan (CK) two decades after devised a general theory of aging. Central in this is the concept of time-reparametrization invariance [22]. The object of study was a spin glass quenched to zero temperature from a high-temperature equilibrium state. Such a system ages continuously and, in fact, never reaches thermal equilibrium. This is in stark contrast to the TN formalism that focuses on relatively small temperature variations [27]. Nevertheless, the CK aging theory is consistent with the existence of a material time (see below). The CK theory is based on the triangular relation [22] that provides a precise definition of time-reparametrization invariance.
The time-autocorrelation function $C_{12}$ of a quantity $A(t)$ is defined by
$C_{12}\equiv\langle A(t_{1})A(t_{2})\rangle-\langle A(t_{1})\rangle\langle A(t_{2})\rangle\,.$ (1)
Here the angular brackets denote an ensemble average, not a time average (recall that time-translational invariance does not apply during aging). We shall assume that all time-autocorrelation functions are positive and go to zero monotonically at large times. The triangular relation states that for all times $t_{1}<t_{2}<t_{3}$ during the aging, any two time-autocorrelation functions determine the third. Specifically,
$C_{13}=F_{A}(C_{12},C_{23})$ (2)
in which $F_{A}(x,y)$ is a function that in general depends on $A$. We note that Eq. (2) applies trivially in thermal equilibrium because by time-translational invariance the time-autocorrelation function $\langle A(t_{1})A(t_{2})\rangle$ is a function of $t_{2}-t_{1}$: $C_{12}$ is a function of $t_{2}-t_{1}$ and $C_{23}$ is a function of $t_{3}-t_{2}$, which implies that $C_{12}$ and $C_{23}$ determine $t_{3}-t_{1}=t_{3}-t_{2}+t_{2}-t_{1}$, which determines $C_{13}$. CK derived Eq. (2) from an approximate mean-field equation for the time development of the time-autocorrelation function. In that equation, if there is a clear time-scale separation between the fast and slow degrees of freedom, then Eq. (2) applies to a good approximation at long times, i.e., for the slow degrees of freedom.
The material time $\xi(t)$ is defined from
$C_{12}=\phi_{\rm eq}(\xi(t_{2})-\xi(t_{1}))\,$ (3)
in which the function $\phi_{\rm eq}$ gives the thermal-equilibrium time-autocorrelation functions via $C_{12}=\phi_{\rm eq}(t_{2}-t_{1})$ at some reference temperature (changing the reference temperature scales $\xi(t)$ by a constant). Equation (2) ensures consistency of this definition by reasoning as above [21]: $\xi(t_{2})-\xi(t_{1})$ determines $C_{12}$ and $\xi(t_{3})-\xi(t_{2})$ determines $C_{23}$, meaning that $\xi(t_{3})-\xi(t_{1})=\xi(t_{3})-\xi(t_{2})+\xi(t_{2})-\xi(t_{1})$ determines $C_{13}$. Each quantity $A$ can have its own material time, but in this paper we exclusively define the material time from the potential energy.
As mentioned, the main characteristic of the TN formalism is the reduction when time is replaced by material time of the generally highly nonlinear aging phenomenon to one described by linear-response theory [8, 2, 21]. The CK approach is very different. Equation (2) is independent of how time is measured, allowing for any smooth monotonic non-linear function of time. This “time-reparametrization invariance” is central in the CK theory of aging, but contrary in philosophy to the TN approach in which the material time plays the special role of linearizing the problem.
Though the definition of the material time $\xi(t)$ in Eq. (3) is straightforward, it only recently became possible to determine this quantity experimentally [28]. This was done from dynamic light scattering data. One reason this has not been done before is that directly measuring a time-autocorrelation function is quite challenging. Moreover, since the statistics cannot be improved by time averaging, one needs to do proper ensemble averaging during the aging. This was implemented in Ref. 28 by analyzing data from the millions of pixels of a camera taking 10-20 pictures per second.
IV Validating the triangular relation in simulations and determining the material time
We first validate the triangular relation Eq. (2). The observed quantity is the potential energy. In order to improve the statistics we consider the potential energies $u_{i}$ of single A particles (defined by allocating half of each interaction energy to each of the two particle involved). The normalized potential-energy time-autocorrelation function is defined by
$C_{uu}(t_{1},t_{2})\equiv\frac{1}{N_{A}\,\sqrt{\nu(t_{1})\,\nu(t_{2})}}\left\langle\sum_{i=1}^{N_{A}}\Delta u_{i}(t_{1})\Delta u_{i}(t_{2})\right\rangle\,.$ (4)
Here $N_{A}$ is the number of A particles, $\Delta u_{i}(t)$ is the potential energy of an A particle minus this quantity’s ensemble average at time $t$ (that is independent of $i$), $\nu(t)\equiv\left\langle\sum_{i=1}^{N_{A}}(\Delta u_{i}(t))^{2}\right\rangle/N_{A}$, and the angular brackets denote an average over 50 simulations. We henceforth leave out the word “normalized” when speaking about time-autocorrelation functions and refer to $C_{uu}(t_{1},t_{2})$ simply as the potential-energy time-autocorrelation function.
The triangular relation is investigated by considering millions of time-autocorrelation functions at times $t_{1}<t_{2}<t_{3}$, evaluated after a temperature up jump carried out at $t=0$. For given values of $C_{12}$ and $C_{23}$ during the aging, Eq. (2) posits that $C_{13}$ is uniquely determined. 250 linearly spaced values for $t_{1}$ were chosen, as well as 60 logarithmically spaced values of $\tau^{\prime}$. Denoting the set of $\tau$ values by $M$, for each value of $t_{1}$ we studied the values $t_{2}=t_{1}+\tau^{\prime}$ for $\tau^{\prime}\in M$, and for each value of $t_{2}$ we studied $t_{3}=t_{2}+\tau^{\prime\prime}$ with $\tau^{\prime\prime}\in M$ (Fig. 2). Of the $250*60^{2}=900,000$ $(t_{1},t_{2},t_{3})$ triplets we kept those for which $t_{3}$ is smaller than the duration of the simulation.
The aging part of the relaxation response occurs at long times. Short times correspond to vibrational relaxation of the system within the cage formed by neighboring particles; here the fluctuation-dissipation theorem applies corresponding to the bath temperature $T$ [22, 29]. We do not show this linear regime since it is of little relevance for the present study. The reader should note, however, that reparametrization by the material time is not expected to collapse the early response (indeed it does not).
To calculate the parametric representation of the triangular relation, $C_{12}$ and $C_{23}$ were each binned into 100 equal sized bins of size 0.05 units of correlation. In this way we obtain a 10,000 pixel grid, and each duplet $\left(C_{12},C_{23}\right)$ falls into a single pixel. As shown in Fig. 2, each pixel contains a distribution of “triangles” with lengths $\left(C_{12},C_{23},C_{13}\right)$. After disregarding pixels with less then 15 triangles, the mean value of $C_{13}$ at each pixel is obtained and showed in Fig. 2. The statistics of this plot is highly sensitive to the sampling time, as well as to the change of the correlation function with $t_{1}$ and $t_{2}$. Ideally, the time sampling should be finest when the autocorrelation changes most such that a dense spectrum of values of the autocorrelation is available. We reconcile this later constraint with the finite storage space available by using a saving scheme composed of many linearly spaced time blocks inside of which saving occurs logarithmically.
Figure 2: Parametric representation of the triangular relation, Eq. (2). (a) Illustration of the sampling of time triplets used to calculate correlation “triangles”. The values on the time axis illustrate the logarithmic increase in time intervals within a single simulation block. (b) Illustration of the triangle binning algorithm. Each pixel contains a distribution of triangles $(C_{12},C_{23},C_{13})$. (c) The average value $\left\langle C_{13}\right\rangle$ as a function of $C_{12}$ for given values of $C_{23}$. This is the mean value used to determine the function $F_{A}$ in Eq. (2). The form of $F_{A}$ averaged over each pixel remains unchanged as the amplitude of the jumps changes, as predicted by the TN formalism in which $F_{A}$ is determined by the equilibrium dynamics. The panel shows that $F_{A}$ is well defined. The standard error $\sigma_{C_{13}}/\sqrt{n}$ is indicated by the width of the colored lines in which $n$ is the number of triangles and $\sigma_{C_{13}}$ the absolute standard deviation of $C_{13}$ in the pixel. (d) shows a heat map of the average $\left\langle C_{13}\right\rangle$ as a function of $C_{12}$ and $C_{23}$.
Figure 3(a) summarizes our findings in the form of heat maps for equilibrium and the two temperature up jumps. The colors indicate the standard deviation defined by $\sigma^{2}\equiv{\langle C_{13}^{2}\rangle-\langle C_{13}\rangle^{2}}$. The left panel shows results for equilibrium simulations in which the triangular relation as mentioned always applies as a consequences of time-translational invariance. The middle and right panels show results for temperature up jumps from the very low temperature 0.37 (at which months of GPU simulations were carried out to attain proper equilibrium configurations [24]). Although there is a slight coloring in the latter case, the fluctuations of $C_{13}$ are minute. Notice that those deviations occur somewhat in the middle of the map corresponding to intermediate values of $t_{2}-t_{1}$.
For a number of temperature jumps, Fig. 3(b) quantifies how far the triangular relation is from being rigorously obeyed. This is done by reporting the number of pixels deviating more than a given threshold. In general the deviations are small, albeit larger than in equilibrium at $T=0.48$ (open black squares). In view of the fact that the triangular relation always applies in equilibrium, the latter data provide a baseline showing the numerical uncertainties.
We conclude that the triangular relation is obeyed to a good approximation for the potential-energy time-autocorrelation functions. As detailed above, this makes it possible to define a material time. Figure 3(c) shows the material time as a function of time for two temperature up jumps to $T=0.48$. At long times equilibrium is approached and the material time becomes proportional to time (the proportionality constant is irrelevant; it depends on the precise definition of $\xi$). At short times the material time develops more slowly with time, an effect that is of course most significant for the large up jump (blue triangles).
Figure 2 plots $C_{13}$ as a function of $C_{12}$ for given values of $C_{23}$. We learn from this plot that the shape of the function $F_{A}$ is the same in equilibrium and for the largest jump, as predicted by the triangular relation. Physically this shows that, despite its strongly nonlinear nature, aging in a sense takes place close to equilibrium [21].
Figure 3: Numerical test of the triangular relation Eq. (2). $\sigma_{C_{13}}$ is the standard deviation of the single-particle potential energy’s time-autocorrelation function $C_{13}\equiv C_{uu}(t_{1},t_{3})$ for fixed values of $C_{12}$ and $C_{23}$. The lower the value of $\sigma_{C_{13}}$ , the better is Eq. (2) satisfied. (a) The absolute standard deviation $\sigma_{C_{13}}$ is close to 0 if Eq. (2) applies. The statistical noise is illustrated by the equilibrium heat map at $T=0.48$ (left panel). In the middle and right heat maps, the deviation is shown for the two temperature up jumps. A small increase in the deviation is observed for the large up jump compared to the smaller one. (b) Number of pixels whose deviation from the triangular relation exceeds a given threshold, plotted as a function of the threshold for a number of different up jumps. A consistent increase of the deviation is observed as the size of the jump is increases, but the largest values of $\sigma_{C_{13}}$ are only roughly twice as large as at equilibrium. The initial number of pixels vary for different jumps because we disregard pixels with too little data. (c) The material time $\xi$ is a non-linear function of time during aging, as shown here for two temperature up jumps. At long times equilibrium at $T=0.48$ is approached and $\xi$ becomes proportional to time.
V Aging following jumps from T=0.43 and T=0.37 to T=0.48
How well does the TN formalism work? Having established that the triangular relation is obeyed for the potential energy, which allows for defining a material time based on this quantity, we proceed to present results for aging of the potential energy and four other quantities following temperature up jumps from well-equilibrated states.
The overall structural relaxation of the system is controlled by the 80% majority particles, the A particles (the B particles are smaller and diffuse faster than the A particles). For this reason we focus entirely on properties relating to the A particles. We studied the aging of the following five quantities (of the A particles):
  1. 1.
    The normalized potential-energy time-autocorrelation function $C_{uu}(t_{1},t_{2})$ defined in Eq. (4).
  2. 2.
    The incoherent intermediate scattering function $F_{s}(t_{1},t_{2})$ calculated from the particle displacement from time $t_{1}$ to time $t_{2}$, evaluated at the wavevector corresponding to the first maximum of the radial distribution function (we determined this at the final temperature $0.48$ but the peak position was virtually constant throughout all simulations),
    $F_{s}(t_{1},t_{2})=\Bigg\langle\frac{1}{N_{A}}\sum_{i=1}^{N_{A}}\cos({\bf q}\cdot({\bf r}_{i}(t_{2})-{\bf r}_{i}(t_{1}))\Bigg\rangle\,.$ (5)
  3. 3.
    The mean-square displacement between times $t_{1}$ and $t_{2}$,
    $\left\langle\Delta r^{2}\right\rangle(t_{1},t_{2})=\Bigg\langle\frac{1}{N_{A}}\sum_{i=1}^{N_{A}}({\bf r}_{i}(t_{2})-{\bf r}_{i}(t_{1}))^{2}\Bigg\rangle\,.$ (6)
  4. 4.
    The dynamic susceptibility $\chi_{4}(t_{1},t_{2})$ quantifying the degree of dynamic heterogeneity in terms of the variance of the incoherent intermediate scattering function [30],
    $\chi_{4}(t_{1},t_{2},q)=N_{A}\Biggl[\Bigg\langle\left(\frac{1}{N_{A}}\sum_{i=1}^{N_{A}}\cos({\bf q}\cdot({\bf r}_{i}(t_{2})-{\bf r}_{i}(t_{1})\right)^{2}\Bigg\rangle-F_{s}^{2}(t_{1},t_{2})\Biggr]\,.$ (7)
    The length of $\bf q$} is the same as above for the incoherent intermediate scattering function;
  5. 5.
    The non-Gaussian parameter $\alpha_{2}(t_{1},t_{2})$ (which is zero for Gaussian diffusion) defined by
    $\alpha_{2}(t_{1},t_{2})=\frac{3\left\langle(\Delta r)^{4}(t_{1},t_{2})\right\rangle}{5\Big(\left\langle(\Delta r)^{2}(t_{1},t_{2})\right\rangle\Big)^{2}}-1\,.$ (8)
Figure 4: Left Two-time functions plotted during aging as functions of the time difference $t_{2}-t_{1}$. The left column gives data for the small up jump, the right for the large one. The color of the curves vary from blue for small values of $t_{1}$ to red as this value increases. The following two-time observables are considered as functions of $t_{2}-t_{1}$: (a-b) Per-particle potential energy autocorrelation function; (c-d) Self part of the intermediate scattering function; (e-f) Mean-square displacement; (g-h) Dynamical susceptibility (Eq. (7)); (i-j) Non-Gaussian parameter (Eq. (8)). The curves are identical at large $t_{1}$ for both jumps because they share the same final temperature while they differ at small $t_{1}$s. Right Same data plotted as functions of the material-time difference $\xi_{2}-\xi_{1}$. There is a significantly better collapse for the small up jump than for the large one.
Figure 4 reports the results of our simulations for the temperature up jumps from $T=0.43$ and $T=0.37$ to $T=0.48$. The left two columns show the time-autocorrelation functions during aging plotted as a function of $t_{2}-t_{1}$ with the smaller jump to the left, parametrized by the waiting time $t_{1}$. As expected, curve collapse is not observed for any of the five quantities. Instead a general speed up is seen, with the short-waiting time bluish curves moving to shorter times. This is the well-known “self-catalyzed” behavior, reflecting the fact that the structure gradually approaches that of equilibrium at $T=0.48$ at which relaxation is much faster than at the two starting temperatures. Comparing the results for the jump from $T=0.43$ with those of the $T=0.37$ to $T=0.48$ jump, there is a considerably larger spread in the latter case. This reflects that fact that the relaxation rates vary more, the larger the difference is between two temperatures. At the longest times corresponding to the red curves, a collapse is observed because equilibrium at $T=0.48$ is approached.
The right two columns of Fig. 4 give the main results of this paper, investigating to which extent the material time collapses all the autocorrelation functions to be unique functions of $\xi_{2}-\xi_{1}$ in which $\xi_{j}$ is the value of the material time at time $t_{j}$. There is no collapse at short material times because, as already mentioned, the material-time concept makes little sense on the short time scale corresponding to phonon times. At larger values of $\xi_{2}-\xi_{1}$ there is a much better collapse, but to a varying degree, and curiously, bluish curves corresponding to short waiting time tend to collapse better then intermediate values of $t_{1}$. Generally, the small temperature jump results in a better collapse than the large one.
Interestingly, the different quantities do not show the same behavior. Starting with the small jump, the potential-energy autocorrelation function and the incoherent intermediate scattering function show quite good collapse, while the mean-square displacement, $\chi_{4}$, and $\alpha_{2}$, do not collapse (though much more so than in the real-time panel). The situation is worse for the large up jump in which case none of the five quantities exhibit a convincing collapse. Even here, however, most of the huge variation of the real-time panel is eliminated by the introduction of the material time. One also observes that while the $\chi_{4}$ curves systematically shift toward later times compared to the final equilibrium value, the $\alpha_{2}$ curves behave oppositely: those at small $t_{1}$ relax faster than the curves at larger $t_{1}$.
VI Discussion
We have seen that the triangular relation is well obeyed for the potential energy. Despite this, the material time defined from the potential energy is not able to describe well the aging of the five autocorrelation functions following, in particular, the largest temperature up jump. Two possible explanations of this comes to mind. One possibility is that each quantity has its own material time in the sense that its aging is much better described in this way than by using a common material time. Another possibility is that dynamic heterogeneities invalidates the existence of any global material time in large up jumps, as suggested some time ago by Castillo and Parsaeian [31]. This would make sense physically because little relaxation can take place in a slow region. Indeed, for a two-dimensional system Herrero et al. have recently demonstrated the existence of growing regions of activity during large temperature up jumps [32]. On the other hand, even equilibrium highly viscous liquids exhibit dynamic heterogeneities, and here there is time-translational invariance of all time-autocorrelation functions and no need for introducing local times. Thus if local clocks are needed to describe large up jumps, this can only be the case under these specific aging conditions. It appears that more work is needed to clarify whether the introduction of local clocks somehow can “save” the TN formalism or it simply has to be discarded for large temperature up jumps.
Acknowledgements.
This work was supported by the VILLUM Foundation’s Matter grant (VIL16515).